William is 1/6 as old as his father but 5 years older than his sister, Mary. If the sum of all their ages is 51, how old is William?
(not from a text) Thank You for your help.

Answer by fractalier(2101)

Let their ages be W, F, and M.
So W = (1/6)F and that means F = 6W.
Also W = M + 5 and that means M = W - 5.
But we are given that W + F + M = 51
Now substitute and get
W + 6W + (W - 5) = 51
8W - 5 = 51
8W = 56
W = 7